a) The temperature of 200 g of water increases from 20°C to 50°C in 10 minutes. Calculate:

   

i. the amount of heat gained by the water.  

   

ii. the average rate at which the water is gaining heat.    

 

  (Specific heat capacity of water = 4200 J kg-1 OC-1)


 

b) Determine the amount of heat that must be removed from 1 kg of water at 40°C to convert it to ice at -5°C.

 


1
Expert's answer
2022-01-16T13:12:39-0500

(a) Given:

"m=\rm 0.2\: kg"

"c=4200\:\rm J/(kg\ ^\circ C)"

"t_i=\rm 20\: ^\circ C"

"t_f=\rm 50\: ^\circ C"

"\tau=\rm 10\: min=600\: s"


(i)

"Q=cm(t_f-t_i)"

"Q=4200*0.2*(50-20)=25200\:\rm J"

(ii)

"P=\frac{Q}{\tau}=\frac{25200}{600}=42\:\rm W"